# Flat quasi-Frobenius Lie superalgebras

### v1

@@ 4. Flat quasi-Frobenius Lie superalgebras
$\mathcal{R}(u,v):={\mathrm{L}}^{\star}_{[u,v]}-[{\mathrm{L}}^{\star}_{u},{\mathrm{L}}^{\star}_{v}],\qquad\forall\,u,v\in\mathfrak{g},$
@@
$\mathcal{R}(u,v)=0,\quad\forall\,u,v\in\mathfrak{g},$
@@
$\omega(\mathrm{L}_{u}^{\star}v,w)=\tfrac{1}{3}\,\omega([u,v],w)+\tfrac{1}{3}(-1)^{|v||w|}\,\omega([u,w],v)=\tfrac{2}{3}\,\omega([u,v],w)+\tfrac{1}{3}(-1)^{|u||v|+|u||w|}\,\omega([v,w],u).$
@@
Since $u\star v=0$, and the fact that ${\mathrm{L}}^{\star}_{u}=-\tfrac{1}{2}{\mathrm{R}}^{\star}_{u}=\tfrac{2}{3}\mathop{\mathrm{ad}}\nolimits_{u}$, for $u\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, by Prop. 4.2.  Proposition, this reduces to
@@
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat orthosymplectic quasi-Frobenius Lie superalgebra. Then the Lie superalgebra $(\mathfrak{g},[\;,\;])$ is solvable.
@@
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat periplectic quasi-Frobenius Lie superalgebra. Then the Lie superalgebra $(\mathfrak{g},[\;,\;])$ is nilpotent.
@@
$\bar{\omega}(\bar{u}\,\bar{\star}\,\bar{v},\bar{w})=\frac{1}{3}\,\bar{\omega}([\bar{u},\bar{v}],\bar{w})+\frac{1}{3}\,\bar{\omega}([\bar{u},\bar{w}],\bar{v}),\qquad\forall\bar{u},\bar{v},\bar{w}\in\bar{\mathfrak{g}}.$
@@
$\omega(U_{i}\star U_{j},V_{k})=\frac{1}{3}(\omega([U_{i},U_{j}],V_{k})+\omega([U_{i},V_{k}],V_{j})$
$=\frac{1}{3}(\sum_{l=1}^{n}C_{ij}^{l}\omega(U_{l},V_{k})+\frac{1}{3}(\sum_{l=1}^{n}E_{ik}^{l}\omega(V_{l},U_{j}))=\frac{1}{3}\sum_{l=1}^{n}C_{ij}^{l}\omega(U_{l},V_{k})-\frac{1}{3}\sum_{l=1}^{n}E_{il}^{j}\omega(U_{l},V_{k})$
@@
We observe that the structure constants $C_{ij}^{k}$ determine the nilpotent Lie algebra structure on $\bar{\mathfrak{g}}_{0}$, while the constants $E_{ij}^{k}$ ensure that the representation $\rho_{\bar{\mathfrak{g}}}$ is nilpotent. Consequently, $\bar{\mathfrak{g}}_{0}$ is nilpotent, and the corresponding adjoint action on $\bar{\mathfrak{g}}_{1}$ is nilpotent as well. This implies that the Lie superalgebra $\mathfrak{g}=\mathfrak{g}_{0}\oplus\mathfrak{g}_{1}$ is nilpotent [RSS]. ∎
@@
Now, suppose $\omega$ is even, and assume that $\mathfrak{g}=[\mathfrak{g},\mathfrak{g}]\oplus[\mathfrak{g},\mathfrak{g}]^{\perp}.$ According to Prop. 4.4.  Theorem, $(\mathfrak{g},[\;,\;])$ is solvable. Hence, $\mathfrak{g},\mathfrak{g}]$ is nilpotent, and its center satisfies $Z([\mathfrak{g},\mathfrak{g}])\neq\{0\}.$
@@
On the other hand, Prop. 4.3.  Proposition implies that $\mathrm{ad}_{v}\circ\mathrm{ad}_{w}=0$, for any $w\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, and hence $[[u,v],w]=0.$ Thus $u,v]\in Z(\mathfrak{g})\subseteq[\mathfrak{g},\mathfrak{g}]^{\perp}$, which implies $[u,v]=0$. This shows that $u\in Z(\mathfrak{g})$.
@@
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat quasi-Frobenius Lie superalgebra. Then $N_{\ell}(\mathfrak{g},\star)$ is a two-sided ideal of $(\mathfrak{g},\star)$ and $\mathfrak{g},[\mathfrak{g},\mathfrak{g}]^{\bot}]\subseteq N_{\ell}(\mathfrak{g},\star)$.
@@ 5.1. Central extension of left-symmetric superalgebras
$(u+\alpha e)\diamond(v+\beta e):=u\bullet v+\mu(u,v)e,\qquad\forall\,u,v\in{\mathcal{A}},\;\alpha,\beta\in\operatorname{\mathbb{K}}.$
@@
(16) $\mu(\mathrm{L}_{u}^{\bullet}(v),w)-\mu(u,\mathrm{L}_{v}^{\bullet}(w))=(-1)^{|u||v|}\big(\mu(\mathrm{L}_{v}^{\bullet}(u),w)-\mu(v,\mathrm{L}_{u}^{\bullet}(w))\big),$
@@ 5.5. Flat double extensions of flat quasi-Frobenius Lie superalgebras
$(u+\alpha e)\diamond(v+\beta e):=u\star v+\omega(\xi(u),v)\,e,\qquad\forall\,u,v\in\mathfrak{b},\;\alpha,\beta\in\operatorname{\mathbb{K}},$
@@
(19) $\xi([u,v]_{\mathfrak{b}})=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),$
for every $u,v\in\mathfrak{b}$.
@@
(20) $\left\{\begin{array}[]{lcllcl}\xi([u,v])&=&\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),&\xi^{*}\circ\xi&=&\tfrac{1}{3}({\mathrm{R}}^{\star}_{b_{0}}+({\mathrm{R}}^{\star}_{b_{0}})^{*}),\\[5.69054pt] D(u\star v)&=&D(u)\star v+u\star D(v)-\xi(u)\star v,&[\xi,\xi^{*}]&=&\xi^{2}-\tfrac{1}{3}{\mathrm{R}}^{\star}_{b_{0}},\;b_{0}\in\ker{D},\end{array}\right$.
@@
$\xi([u,v])=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),\quad\text{and}\quad\xi^{*}\circ\xi=\tfrac{1}{3}({\mathrm{R}}^{\star}_{b_{0}}+({\mathrm{R}}^{\star}_{b_{0}})^{*}).$
@@
$(u+\alpha e)\star(v+\beta e)=u\star_{\mathfrak{b}}v+\omega_{\mathfrak{b}}(\xi(u),v)e,$
for all elements $u,v\in\mathfrak{b}$ and $\alpha,\beta\in\mathbb{K}$. By Lemma 5.5.2.  Lemma, the first condition of system (20), for all $u,v\in\mathfrak{b}$
@@
Let $a\in\mathfrak{g}_{\bar{0}}$. Then, we have $\omega({\mathrm{L}}^{\star}_{a}(u),v)=-\,\omega(u,{\mathrm{L}}^{\star}_{a}(v)),$ for all $u,v\in\mathfrak{g}$. Choose a homogeneous basis (even $\mid$ odd) $\{X_{1},Y_{1},\ldots,X_{n},Y_{n},\mid f_{1},\ldots,f_{m}\}$ of $\mathfrak{g}$ such that
@@
$={\mathrm{R}}^{\star}_{a}\circ\Big(({\mathrm{R}}^{\star}_{a})^{\,n+1}+({\mathrm{R}}^{\star}_{a})^{\,n-2}\circ[{\mathrm{L}}^{\star}_{a},{\mathrm{R}}^{\star}_{a}]+({\mathrm{R}}^{\star}_{a})^{\,n-3}\circ[{\mathrm{L}}^{\star}_{x_{1}},{\mathrm{R}}^{\star}_{a}]+\dots+[{\mathrm{L}}^{\star}_{x_{n-1}},{\mathrm{R}}^{\star}_{a}]\Big)+[{\mathrm{L}}^{\star}_{x_{n}},{\mathrm{R}}^{\star}_{a}]$
@@
$\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k}\circ[{\mathrm{L}}^{\star}_{a},{\mathrm{R}}^{\star}_{a}])$ $=\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k}\circ({\mathrm{L}}^{\star}_{a}\circ{\mathrm{R}}^{\star}_{a}-{\mathrm{R}}^{\star}_{a}\circ{\mathrm{L}}^{\star}_{a})=\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k}\circ({\mathrm{L}}^{\star}_{a}\circ{\mathrm{R}}^{\star}_{a}))-\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k+1}\circ{\mathrm{L}}^{\star}_{a}))$

### v2

@@ 4. Flat quasi-Frobenius Lie superalgebras
$\mathcal{R}(u,v):={\mathrm{L}}^{\star}_{[u,v]}-[{\mathrm{L}}^{\star}_{u},{\mathrm{L}}^{\star}_{v}],\qquad\text{ for all }u,v\in\mathfrak{g},$
@@
$\mathcal{R}(u,v)=0,\quad\text{ for all }u,v\in\mathfrak{g},$
@@
Example.
In this example, we study an important class of nilpotent Lie superalgebras, known as filiform Lie superalgebras. We restrict our attention to the simplest ones, denoted by $L^{n,m}$, where $n,m\in\mathbb{N}$. These are defined on the homogeneous basis
$(e_{1},\dots,e_{n}\mid f_{1},\dots,f_{m}),$
whose only nonzero bracket products are given by
$[e_{1},e_{i}]$ $=e_{i+1},\qquad 2\leq i\leq n-1,$
$[e_{1},f_{j}]$ $=f_{j+1},\qquad 1\leq j\leq m-1.$
The case where $m=n$. Following [BM], $L^{n,n}$ admits a closed periplectic form given by
$\omega=\lambda\,e_{1}^{*}\wedge f_{n}^{*}+\mu\sum_{i=2}^{n}(-1)^{\,n-i}e_{i}^{*}\wedge f_{\,n+1-i}^{*},\quad\text{where $\lambda\mu\neq 0$.}$
Using Eq (13), the associated natural symplectic product $\star$ is given by: (non-displayed products are zero)
$e_{1}\star e_{i}$ $=\frac{2}{3}e_{i+1},$ $\qquad e_{i}\star e_{1}$ $=-\frac{1}{3}e_{i+1},\quad e_{i}\star f_{n-i}=f_{n-i}\star e_{i}=\frac{\mu}{3\lambda}(-1)^{n+1-i}f_{n},$
$e_{1}\star f_{j}$ $=\frac{2}{3}f_{j+1},$ $\qquad f_{j}\star e_{1}$ $=-\frac{1}{3}f_{j+1},\quad e_{1}\star e_{1}=\frac{\lambda}{3\mu}(-1)^{n+1}e_{2},$
$e_{1}\star f_{n-1}$ $=\frac{1}{3}f_{n},$ $\qquad f_{n-1}\star e_{1}$ $=-\frac{2}{3}f_{n},$
where $2\leq i\leq n-1$ and $1\leq j\leq n-2$.
A direct computation gives
$\mathcal{R}(e_{1},e_{i})(e_{1})=\big({\mathrm{L}}^{\star}_{[e_{1},e_{i}]}-[{\mathrm{L}}^{\star}_{e_{1}},{\mathrm{L}}^{\star}_{e_{i}}]\big)(e_{1})=-\frac{1}{9}\,e_{i+2}\neq 0.$
Therefore, $(L^{n,n},[\;,\;],\omega)$ is not flat.
The case where $n$ is even and $m$ is odd. Following [BM], the Lie superalgebra $L^{n,m}$ admits a closed orthosymplectic form given by
$\omega=\lambda\,e_{1}^{*}\wedge e_{n}^{*}+\mu\sum_{i=2}^{\frac{n}{2}}(-1)^{\frac{n}{2}-i}e_{i}^{*}\wedge e_{n+1-i}^{*}+(-1)^{\frac{m+1}{2}}\nu\sum_{i=1}^{\frac{m-1}{2}}(-1)^{i}f_{i}^{*}\wedge f_{m+1-i}^{*}+\frac{\nu}{2}f_{\frac{m+1}{2}}^{*}\wedge f_{\frac{m+1}{2}}^{*},$
where $\lambda\mu\nu\neq 0$.
Using again Eq. (13), one obtains the associated product $\star$ as in the previous case. In particular, $(L^{n,m},[\;,\;],\omega)$ is not flat. We omit the details.
The case where $n$ is even and m=n2m=n-2. Following [BM], the Lie superalgebra $L^{n,n-2}$ admits a closed, antisymmetric, nondegenerate form that is nonhomogeneous. This lies beyond the scope of the present paper.
The case where $n$ is even and $m\not=n-2$. It was conjectured in [BM] that no closed and nondegenerate form exists.
@@
$\begin{array}[]{lcl}\omega(\mathrm{L}_{u}^{\star}v,w)&=&\tfrac{1}{3}\,\omega([u,v],w)+\tfrac{1}{3}(-1)^{|v||w|}\,\omega([u,w],v)\\[5.69054pt] &=&\tfrac{2}{3}\,\omega([u,v],w)+\tfrac{1}{3}(-1)^{|u|(|v|+|w|)}\,\omega([v,w],u).\end{array}$
@@
Since $u\star v=0$, and the fact that ${\mathrm{L}}^{\star}_{u}=-\tfrac{1}{2}{\mathrm{R}}^{\star}_{u}=\tfrac{2}{3}\mathop{\mathrm{ad}}\nolimits_{u}$, for $u\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, by Prop. 4.3.  Proposition, this reduces to
@@
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat orthosymplectic quasi-Frobenius Lie superalgebra. Then, the Lie superalgebra $(\mathfrak{g},[\;,\;])$ is solvable.
@@
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat periplectic quasi-Frobenius Lie superalgebra. Then, the Lie superalgebra $(\mathfrak{g},[\;,\;])$ is nilpotent.
@@
$\bar{\omega}(\bar{u}\,\bar{\star}\,\bar{v},\bar{w})=\frac{1}{3}\,\bar{\omega}([\bar{u},\bar{v}],\bar{w})+\frac{1}{3}\,\bar{\omega}([\bar{u},\bar{w}],\bar{v}),\qquad\text{for all }\bar{u},\bar{v},\bar{w}\in\bar{\mathfrak{g}}.$
@@
$\omega(U_{i}\star U_{j},V_{k})=\frac{1}{3}\omega([U_{i},U_{j}],V_{k})+\frac{1}{3}\omega([U_{i},V_{k}],V_{j})$
$=\frac{1}{3}\sum_{l=1}^{n}C_{ij}^{l}\omega(U_{l},V_{k})+\frac{1}{3}\sum_{l=1}^{n}E_{ik}^{l}\omega(V_{l},U_{j})=\frac{1}{3}\sum_{l=1}^{n}C_{ij}^{l}\omega(U_{l},V_{k})-\frac{1}{3}\sum_{l=1}^{n}E_{il}^{j}\omega(U_{l},V_{k})$
@@
We observe that the structure constants $C_{ij}^{k}$ determine the nilpotent Lie algebra structure on $\bar{\mathfrak{g}}_{0}$, while the constants $E_{ij}^{k}$ ensure that the representation $\rho_{\bar{\mathfrak{g}}}$ is nilpotent. Consequently, $\bar{\mathfrak{g}}_{0}$ is nilpotent, and the corresponding adjoint action on $\bar{\mathfrak{g}}_{1}$ is nilpotent as well. This implies that the Lie superalgebra $(\mathfrak{g}=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{\bar{1}},[\;,\;])$ is nilpotent [RSS]. ∎
Remark.
The reader may wonder why the argument in Theorem 4.6.  Theorem fails in the orthosymplectic case. The point is that, upon forgetting the superstructure, the Lie algebra $(\bar{{\mathfrak{g}}},[\;,\;],\bar{\omega})$ need not be symplectic. Therefore, the result of [BELL] does not apply, and one cannot deduce that $(\bar{{\mathfrak{g}}},[\;,\;])$ is nilpotent.
@@
Now, suppose $\omega$ is even, and assume that $\mathfrak{g}=[\mathfrak{g},\mathfrak{g}]\oplus[\mathfrak{g},\mathfrak{g}]^{\perp}.$ According to Prop. 4.5.  Theorem, $(\mathfrak{g},[\;,\;])$ is solvable. Hence, $\mathfrak{g},\mathfrak{g}]$ is nilpotent, and its center satisfies $Z([\mathfrak{g},\mathfrak{g}])\neq\{0\}.$
@@
On the other hand, Prop. 4.4.  Proposition implies that $\mathrm{ad}_{v}\circ\mathrm{ad}_{w}=0$, for any $w\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, and hence $[[u,v],w]=0.$ Thus, $u,v]\in Z(\mathfrak{g})\subseteq[\mathfrak{g},\mathfrak{g}]^{\perp}$, which implies $[u,v]=0$. This shows that $u\in Z(\mathfrak{g})$.
@@
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat quasi-Frobenius Lie superalgebra. Then, $N_{\ell}(\mathfrak{g},\star)$ is a two-sided ideal of $(\mathfrak{g},\star)$ and $\mathfrak{g},[\mathfrak{g},\mathfrak{g}]^{\bot}]\subseteq N_{\ell}(\mathfrak{g},\star)$.
@@ 5.1. Central extension of left-symmetric superalgebras
$(u+\alpha e)\diamond(v+\beta e):=u\bullet v+\mu(u,v)e,\qquad\text{for all }u,v\in{\mathcal{A}},\;\alpha,\beta\in\operatorname{\mathbb{K}}.$
@@
(16) $\mu(\mathrm{L}_{u}^{\bullet}(v),w)-\mu(u,\mathrm{L}_{v}^{\bullet}(w))=(-1)^{|u||v|}\big(\mu(\mathrm{L}_{v}^{\bullet}(u),w)-\mu(v,\mathrm{L}_{u}^{\bullet}(w))\big),\quad\text{for all $u,v,w\in{\mathcal{A}}$}.$
@@ 5.5. Flat double extensions of flat quasi-Frobenius Lie superalgebras
We explain how the double extension is constructed in the following diagram.
$({\mathfrak{b}},[\;,\;]_{{\mathfrak{b}}},\omega_{\mathfrak{b}})$$({\mathfrak{b}},\bullet)$$({\mathfrak{b}}\oplus V^{*},\diamond)$$({\mathfrak{g}}:=V\oplus{\mathfrak{b}}\oplus V^{*},{\bar{\bullet}})$$({\mathfrak{g}},[\;,\;],\omega)$}Step 1Step 2Step 3Step 4
In the above diagram, we now describe the construction in detail.
Step 1. Select the natural symplectic product associated with $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$.
Step 2. Construct the central extension of $(\mathfrak{b},\bullet)$ (see Section 5.1), where the cocycle is given by
$\mu(\cdot,\cdot)=\omega_{\mathfrak{b}}(\xi(\cdot),\cdot),$
and $\xi$ satisfies the conditions ensuring that equation (16) holds.
Step 3. Form the semidirect product as in Section 5.2, using the data
$(D=\xi^{*}-\xi,\xi,b_{\bar{0}},\lambda),$
subject to the conditions that guaranty that equation (17) is valid.
Step 4. Compute the commutator $[\;,\;]_{\bar{\bullet}}$ to obtain the Lie bracket on $\mathfrak{g}$, and extend the form $\omega_{\mathfrak{b}}$ naturally to $\mathfrak{g}$.
@@
$(u+\alpha e)\diamond(v+\beta e):=u\star v+\omega(\xi(u),v)\,e,\qquad\text{for all }u,v\in\mathfrak{b},\;\alpha,\beta\in\operatorname{\mathbb{K}},$
@@
(19) $\xi([u,v]_{\mathfrak{b}})=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),\quad\text{for all $u,v\in\mathfrak{b}$.}$
@@
(20) $\left\{\begin{array}[]{lcllcl}\xi([u,v]_{\mathfrak{b}})&=&\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),&\xi^{*}\circ\xi&=&\tfrac{1}{3}({\mathrm{R}}^{\star}_{b_{0}}+({\mathrm{R}}^{\star}_{b_{0}})^{*}),\\[5.69054pt] D(u\star v)&=&D(u)\star v+u\star D(v)-\xi(u)\star v,&[\xi,\xi^{*}]&=&\xi^{2}-\tfrac{1}{3}{\mathrm{R}}^{\star}_{b_{0}},\;b_{0}\in\ker{D},\end{array}\right$.
@@
$\widetilde{\xi}([u,v]_{\mathfrak{b}})=\mathrm{L}_{u}^{\diamond}(\widetilde{\xi}(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\diamond}(\widetilde{\xi}(u))$
@@
$\xi([u,v]_{\mathfrak{b}})=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),\quad\text{and}\quad\xi^{*}\circ\xi=\tfrac{1}{3}({\mathrm{R}}^{\star}_{b_{0}}+({\mathrm{R}}^{\star}_{b_{0}})^{*}).$
@@
$(u+\alpha e)\star(v+\beta e)=u\star_{\mathfrak{b}}v+\omega_{\mathfrak{b}}(\xi(u),v)e,\quad\text{for all $u,v\in\mathfrak{b}$ and $\alpha,\beta\in\mathbb{K}$.}$
By Lemma 5.5.2.  Lemma, the first condition of system (20), for all $u,v\in\mathfrak{b}$
@@
Let $a\in\mathfrak{g}_{\bar{0}}$. Then, we have $\omega({\mathrm{L}}^{\star}_{a}(u),v)=-\,\omega(u,{\mathrm{L}}^{\star}_{a}(v)),$ for all $u,v\in\mathfrak{g}$. Choose a homogeneous basis (even $\mid$ odd) $\{X_{1},Y_{1},\ldots,X_{n},Y_{n}\mid f_{1},\ldots,f_{m}\}$ of $\mathfrak{g}$ such that
@@
$={\mathrm{R}}^{\star}_{a}\circ\Big(({\mathrm{R}}^{\star}_{a})^{\,n+1}+({\mathrm{R}}^{\star}_{a})^{\,n-2}\circ[{\mathrm{L}}^{\star}_{a},{\mathrm{R}}^{\star}_{a}]+({\mathrm{R}}^{\star}_{a})^{\,n-3}\circ[{\mathrm{L}}^{\star}_{x_{1}},{\mathrm{R}}^{\star}_{a}]+\dots+[{\mathrm{L}}^{\star}_{x_{n-1}},{\mathrm{R}}^{\star}_{a}]\Big)$
$\quad+[{\mathrm{L}}^{\star}_{x_{n}},{\mathrm{R}}^{\star}_{a}]$
@@
$\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k}\circ[{\mathrm{L}}^{\star}_{a},{\mathrm{R}}^{\star}_{a}])$ $=\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k}\circ({\mathrm{L}}^{\star}_{a}\circ{\mathrm{R}}^{\star}_{a}-{\mathrm{R}}^{\star}_{a}\circ{\mathrm{L}}^{\star}_{a}))$
$=\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k}\circ({\mathrm{L}}^{\star}_{a}\circ{\mathrm{R}}^{\star}_{a}))-\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k+1}\circ{\mathrm{L}}^{\star}_{a}))$